OpenStudy (anonymous):

Suppose Wal-Mart stock is selling at $59 a share and Mattel stock is selling at $28 a share. Amy has a maximum of $6000 to invest. She wishes to purchase four times as many shares of Wal-Mart as of Mattel. Only whole shares of stock can be purchased. a) How many shares of each will she purchase? b) How much money will be left over?

OpenStudy (anonymous):

264x=6000 x= 6000/264 x= 22 shares of Mattel 4x= 4*22= 88 shares of Wal-Mart 88*59+22*28= $ 5808 $ 6000-$5808=$192 balance